OpenStudy (anonymous):

Please Help!! Find an equation for the nth term of a geometric sequence where the second and fifth terms are -21 and 567, respectively.

OpenStudy (anonymous):

@SolomonZelman @precal @paki @jdoe0001

OpenStudy (precal):

well you need to use your formulas for geometric sequences to figure out your common ratio

OpenStudy (solomonzelman):

\(\large\color{blue}{ a_2 \times r^{5-2}=a_5 }\)

OpenStudy (anonymous):

-12*r^3=567

OpenStudy (anonymous):

-21*

OpenStudy (anonymous):

r^3=588

OpenStudy (jdoe0001):

to get from the 2nd term to the 5th term you'd need to multiply the currrent term TIMES some number, namely the "common ratio" or "r" so as SolomonZelman said, it went from the 2nd to 5th, that 3 terms furhther up meaning "r" was used 3 times, so \(\bf r^3\)

OpenStudy (anonymous):

576/-21

OpenStudy (solomonzelman):

\(\large\color{blue}{ a_2 \times r^{5-2}=a_5 }\) \(\large\color{blue}{ (-12) \times r^{3}=567 }\) \(\large\color{blue}{ r^{3}=567 \div (-12)}\) \(\large\color{blue}{ r^{3}=-47.25}\)

OpenStudy (solomonzelman):

Oh it is 21, not 12. Sorry.... but you get the idea.

OpenStudy (anonymous):

-27=r^3 7=r

OpenStudy (solomonzelman):

\(\large\color{blue}{ -27= (-3)^3}\)

OpenStudy (solomonzelman):

\(\large\color{blue}{ 27= 3^3}\)

OpenStudy (jdoe0001):

hmm \(\Large \bf -21r^3=567\implies r=\sqrt[3]{-\cfrac{\cancel{ 567 }}{\cancel{ 21 }}}\)

OpenStudy (anonymous):

so its a_n=7*3^n-1

OpenStudy (solomonzelman):

Yes

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thanks!!

OpenStudy (jdoe0001):

yw

OpenStudy (solomonzelman):

`\(\large\color{black}{ a_n = 7 \times 3 ^ {n -1 } }\) ` Copy and paste the text:) @ninjasandtigers

OpenStudy (anonymous):

Is it Solomon Zelalem

OpenStudy (solomonzelman):

No, close:) Zelman

OpenStudy (solomonzelman):

Have a good luck !